OpenStudy (anonymous):

27x^3+64y^3 will medal

OpenStudy (anonymous):

@FibonacciChick666 please help

OpenStudy (anonymous):

srry i dont like you so i cant help

OpenStudy (anonymous):

krystal

OpenStudy (anonymous):

dont care

OpenStudy (anonymous):

bye i dont want you to help @stephaniehoran23 anyway you are my sister

OpenStudy (thomas5267):

What is the question?

OpenStudy (anonymous):

27x^3+64y^3 i need this factored if you can help i will medal you

OpenStudy (anonymous):

@Kainui @inowalst @ikram002p @ganeshie8 @bohotness can anybody help?

OpenStudy (thomas5267):

\[ a^3+b^3=(a+b)(a^2-ab+b^2) \]

OpenStudy (anonymous):

ok so how do i get the answer? i still do not understand.

OpenStudy (anonymous):

\[27x^3+64y^3\] please help i need to know how to do this i will medal if helped

OpenStudy (anonymous):

it needs to be factored

OpenStudy (thomas5267):

Note that \(27x^3=(3x)^3\). Try to the same thing on the other term.

OpenStudy (anonymous):

so \[64y^3=(4y)^3\] right?